The pressure and temperature of mixture of 4 kg of O2 and 6 kg of N2 are 4 bar 

and 27°C respectively. For the mixture determine the following ) :i) The mole 

fraction of each component ; (ii) The average molecular weight; (iii) The specific 

gas constant ; (iv) The volume and density ;(v) The partial pressures and partial 

volumes


1
Expert's answer
2021-07-12T02:12:02-0400

Mole fraction of oxygen

"n_{o_2}=\frac{gm weight}{moleculer weight}=\frac{4}{32}=0.125"

"n_{N_2}=\frac{gm weight}{moleculer weight}=\frac{6}{28}=0.214"

Molecular weight

Molecular weight for

oxygen =2"\times" atomic weights

"(W_m)_{o_2}=2\times16=32"

Nitrogen molecular weight

="2\times" atomic weights

"(W_m)_{N_2}=2\times14=28"

Specific gas constan

"R_s=\frac{R}{M{_o}_{2}}=\frac{8.31}{32}=0.2596"

"R_s=\frac{R}{M{_N}_{2}}=\frac{8.31}{28}=0.296"

Volume and density

"V=\frac{nRT}{P}"

"V_o=\frac{0.125\times8.314\times300}{4}=77.94m^3"

"V_N=\frac{0.214\times8.314\times300}{4}=133.43m^3"

"d_o=\frac{M}{V}=\frac{32}{77.94}=0.4105kg/m^3"

"d_o=\frac{M}{V}=\frac{28}{133.43}=0.2098kg/m^3"

Partial pressure

"P_o=n\times" pressure

"P_o=0.125\times4=0.5 bar"

"P_N=n\times" pressure

"P_N=" "0.214\times4=0.856" bar

Partial volume

"V_o=n\times" volume

"V_o=0.125\times77.94=9.74kg"

"V_N=n\times" volume

"V_N=0.214\times133.43=28.55kg"